This implies that the production function for long-distance telephone service post-merger will be given by Q = x [K + 1/2L], where x > 1 is the productivity factor. Suppose that r = 10 and w = 5, and the pre-merger service quality index is s = 12.

a) Suppose that the long-distance market is a monopoly following the merger between AT&T and Sprint and that x = 5/4. Assume that s remains constant at 12. Should the DOJ allow this merger to proceed? Provide a careful economic analysis in support of your recommendation.(20)

b) Continue to assume that x = 5/4, but the market for long-distance telephone service remains perfectly competitive following the merger. For what values of s will the DOJ approve this merger? (15)
1
Expert's answer
2020-06-01T13:06:25-0400

(a)


"Q=x[K+\frac{1}{2}L]"


"Q=\frac{5}{4}\times12.5"


"Q=15.625"

"Q=a-b\times p+s"

"15.625=a-b \times p+12"


"bp=a-3.625"


"p=\frac{a-3.625}{b}"

The merger will be allowed in the event that the price of final services does not change.


(b)


"15.625=a-bp+s"


"s\le 15.625 - a +bp"